OpenStudy (anonymous):

********************help please******************** solve for x: |2x+6|-4=20

OpenStudy (anonymous):

What happens when you add 4 to each side?

OpenStudy (johnweldon1993):

Right, we add 4 to both sides \[\large |2x + 6| = 24\] And we break it into 2 equations \[\large 2x + 6 = 24\] and \[\large 2x + 6 = -24\]

OpenStudy (anonymous):

|2x+6|-4=20 |2x+6| =24 2x+6 = 24 ; 2x+6=-24 2x = 18 ; 2x = -30 x = 9 ; x = -15

OpenStudy (muzzack):

yes just like @johnweldon1993 now we do the equations 2x=18 and 2x=-30 then x=9 and x=-15